Researchers suspect that 18% of all high school students smoke at least one pack of cigarettes a day. At Wilson High School, with an enrollment of 300 students, a study found that 50 students smoked at least one pack of cigarettes a day. At a=0.05, test the claim that 18% of all high school students smoke at least one pack of cigarettes day.


1
Expert's answer
2021-12-06T13:22:34-0500

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\geq 0.18"

"H_1:p<0.18"

This corresponds to a left-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a left-tailed test is "z_c = -1.6449."

The rejection region for this left-tailed test is "R = \{z: z < -1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}"

"=\dfrac{50/300-0.18}{\sqrt{\dfrac{0.18(1-0.18)}{300}}}\approx-0.601113"

Since it is observed that "z = -0.601113 \ge -1.6449=z_c ," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value is "p=P(Z<-0.601113)""=0.273882," and since "p = 0.273882>0.05=\alpha ," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion  "p" is less than "0.18," at the "\alpha = 0.05"  significance level.

Therefore, there is enough evidence to claim that at least 18% of all high school students smoke at least one pack of cigarettes day at the "\alpha = 0.05"  significance level.